OpenStudy (anonymous):

Solve the rational equation. 5/x - 1−3/x + 2= 0

OpenStudy (anonymous):

by taking LCM=x we get 5-3-x+2x/x=0 cross multiplying 2+x=0 x=-2 -2 is the answer